A stadium has 49000 seats. Seats sell for $25 in Section A, $20 in Section B, and $15 in Section C. The number of seats in Section A equals the total number of seats in Section B and C. Suppose the stadium takes in $1052000 from each sold-out event. How many seats does each section hold?
Answer by mananth(16946)

A stadium has 49000 seats.
Seats sell for $25 in Section A, z
$20 in Section B,-------------------x seats
$15 in Section C. ------------------y seats
(x+y)=z
25(x+y)+20x+15y=1052000
25x+25y+20x+15y=1052000
45x+40y=1052000
/5
9x+8y=210400------------------1
2x+2y=49000
/2
x+y=24500
multiply by -9
-9x-9y=-220500----------------2
add equation (1) & (2)
-y=-10100
y= 10100 Section C
plug value of y in equation x+y = 24500
x+10100=24500
x=14400 seats Section B
x+y = Z= 24500 Section A
CHECK
24500+14400+10100= 49,000
24500*25+14400*20+15*10100=1,052,000
m.ananth@hotmail.ca